OpenStudy (anonymous):

A and B are the points on the curve y=2^x with the coordinates 3 and 3.1 respectively. Find the gradient/slope of the chord AB and stating the points you use, find the gradient of another chord which will give a closer approx. to the gradient/slope of the tangent to y=2^x at A. thanks

OpenStudy (amistre64):

A = (3, 2^3) = (3,8) B = (3.1,2^(3.1)) = (3.1, 8.57) slope = .57/.1 = 5.7

OpenStudy (amistre64):

since a is at 3; use a closer value of x that is between 3 and 3.1; such as 3.01

OpenStudy (amistre64):

as you get closer and closer to A, which is (3,8); you will approach the limit that is equal to the slope at 'A'; also known as the derivative of y = 2^x at x=3

OpenStudy (amistre64):

does that help?

OpenStudy (anonymous):

wait let me read it through a few times :S

OpenStudy (anonymous):

how did you get the slope? where did you get the numbers from 0.57/0.1=5.7???

OpenStudy (amistre64):

you are given a function for y: 2^x correct? At point A it says that x=3 and at point B, x=3.1 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

point A = (3,2^3) point B = (3.1, 2^(3.1))

OpenStudy (amistre64):

B-A gives you the slope: x , y -------------- 3.1 , 2^(3.1) -3 , - 2^3 --------------- ( .1 , .57418...) slope = y/x = .57/.1 = 5.7/1 = 5.7

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome :)